State Bonnet's mean value theorem for integrals.

Apply it to show that:

5

| ∫ (cos x/x)dx | ≤ 2/3

3


1
Expert's answer
2021-05-19T08:29:27-0400

BONNET MEAN VALUE THEOREM


If both f:[a,b] "\to" R and g:[a,b]"\to" R Integrable on [a,b] and f is non negative and monotonically decreasing on [a,b] then there exist a point "\gamma" "\isin" [a,b] such that "\int_a^b f(x)g(x)dx=f(a)\int_a^\gamma g(x)dx"


Now we need to show that | "\int_3^5 (cosx/x) dx| \leq \frac 2 3"


Here f(x)="\frac 1 x" & g(x) = "cosx"

Here both f(x) and g(x) are integrable on [3,5] and f(x) is non negative and monotonic decreasing on [3,5] .So bonnet theorom is applicable here . Here f(a)=f(3)="\frac 1 3"

So "\int _3^5 (cosx/x)dx = \frac 1 3 \int _3^\gamma (cosx) dx"

="\frac 1 3 [sin(\gamma)-sin(3)]"


Now | "\int_3^5 (cosx/x) dx | = |\frac 1 3 (sin(\gamma)-sin(3))|"

"\leq \frac 1 3 ( |sin(\gamma)|+|sin(3)|)"

"\leq\frac 1 3 (1 +1)"

"\leq\frac 2 3" (Proved)